Suppose you are working with a data set that is normally distributed with a mean of 400 and a standard deviation of 20. Determine the value of x such that 5% of the values are less than x.
Answer by Boreal(15235)

The 5% part of the normal curve is at z=-1.645
z=(x-mean)sd=(x-400)/20
-32.9=x-400
x=367.1
calculator 2nd VARS invnorm (.05,400,20)